# ISRE-0001-2024.00

### v1

@@
[…]
ISRE-0001-2024.00
[…]
Li, De-Arteaga, and Saar-Tsechansky
[…]
Bias-Aware Mislabeling Detection via Decoupled Confident Learning
[…]
Bias-Aware Mislabeling Detection via Decoupled Confident Learning
[…]
@@
[…]
Data integrity, Data quality, Bias-aware design science, Human AI interaction
@@ 1 Introduction
In an era defined by rapid digitization and fueled by advances in computational technologies and algorithms, organizations are increasingly reliant on data to drive competitiveness, inform strategy, and support core operations. Evidence from industry underscores the transformative nature of data-driven practices, as data-driven organizations outperform peers by a factor of 23 in customer acquisition, achieve ninefold improvements in customer retention, and realize up to 19 times greater profitability (McKinsey 2014, Davenport et al. 2006).
@@
Simultaneously, as the strategic value of data-driven practices has intensified, the cost of poor data quality has emerged as a critical point of vulnerability. Analytical systems built on systematic errors in data trigger adverse downstream consequences—producing misleading insights and information that can distort managerial decisions, risk operational disruptions and compliance failures, and erode strategic value and organizational trust (Ballou and Pazer 1985, Bai et al. 2012, Krishnan et al. 2005, Hazen et al. 2014, Parssian et al. 2004). The estimated financial implications of poor data quality are profound, with costs reaching 8% to 12% of revenues for typical organizations, potentially leading to billions of dollars in annual losses (Hazen et al. 2014). Data integrity is at the core of contemporary business competitiveness, which increasingly relies on systematic data auditing and relabeling to enhance data quality (Ganju et al. 2022, Liu 2024, Bernhardt et al. 2022).
@@
A particularly consequential data integrity challenge stems from the quality of labels (Frénay and Verleysen 2013). A Label is a covariate that is typically the key focus of an analysis, the result of which often informs impactful decisions and policies at scale and advances knowledge discovery. For instance, a medical diagnosis functions as a label that reflects the presence and severity of a condition and is fundamental to inform treatment decisions. Similarly, in the context of content moderation, a post may be labeled as ‘toxic’ if it contains hate speech or derogatory content (Davani et al. 2023), based on which the platform may intervene to mitigate risks. The term “label” has been coined and used extensively in the machine learning literature, while different terms for this concept—such as dependent variable, response variable, outcome, annotation, or rating—exist across disciplines, including social sciences, where these terms similarly refer to an outcome of consequence, central to the motivation and implications of the analyses.
Label bias is a particular form of label error that has attracted significant attention across research communities, including Information Systems (IS) and computer science, because it has been shown to be both common and consequential across domains (Obermeyer et al. 2019, Fu et al. 2020, Bjarnadóttir and Anderson 2020, Kokkodis and Ipeirotis 2021, Fu et al. 2022, Li et al. 2024, Jacobs and Wallach 2021, Passi and Barocas 2019, Suresh and Guttag 2021, Sap et al. 2019, Davani et al. 2023, ProPublica 2016, Zanger-Tishler et al. 2024, Hunter et al. 1979, Sjoding et al. 2020, Fogliato et al. 2020, Akpinar et al. 2021).
@@
Label bias refers to systematic errors in the observed labels relative to the gold standard labels, such that the quality of labels differs across social groups (Li et al. 2022, Zanger-Tishler et al. 2024). Such bias can result in disproportionate harms, with some groups being more adversely affected than others. Specifically, under this bias, the likelihood that an instance is incorrectly labeled depends not only on its true label—a dependency that determines what type of error is more common—but also on its group membership (e.g., age bracket, race, gender, or socioeconomic status). A dependency on both true label and group membership means that the types of errors as well as the rates of errors may vary across groups. Label bias commonly arises in settings where the true label is unobservable or costly to obtain, and observed labels are instead derived from proxies, human judgments or measurement instruments that can embed label bias due to social structural biases, cognitive heuristics, or gaps in domain knowledge (Li et al. 2022, 2024, Passi and Barocas 2019, Jacobs and Wallach 2021, Eickhoff 2018, Draws et al. 2021).
@@
Prior research has revealed the adverse impact of label bias across systems and analyses that utilize the labels, demonstrating that such errors undermine productive data-driven decisions across domains, and potentially perpetuate existing inequalities (Li et al. 2024, 2022, Sap et al. 2019, Davani et al. 2023). Analyses that are crucially impacted by labeling quality include data-driven hypothesis generation and pattern discovery, as well as predictive modeling, descriptive modeling, and prescriptive modeling that inform understandings and policies (Agarwal and Dhar 2014, Shmueli 2010, Abbasi et al. 2024). These risks have also been echoed by federal agencies and policies, including by the National Institute of Standards and Technology (NIST), which urged organizations to enhance data quality  (NIST 2023).
@@
For example, health care systems often aim to estimate health needs to inform the allocation of health resources. However, health needs are often prohibitively costly to estimate. Consequently, health care costs—which are abundant and readily available in the U.S.—have been used as a proxy label for health needs (Obermeyer et al. 2019, Li et al. 2022). However, there is a systematic mismatch between health costs and actual health needs, and this mismatch varies across groups. Specifically, because of historical inequities in access to care across patient groups, Black patients have often incurred lower health care costs than white patients with comparable medical conditions. Consequently, a system that relied on cost as a proxy for needs systematically underestimated Black patients’ needs (Obermeyer et al. 2019). Label quality has been impacted by label bias across a range of other contexts, including criminal justice (ProPublica 2016, Zanger-Tishler et al. 2024), content moderation Davani et al. (2023, 2022), Sap et al. (2019), and hiring (Hunter et al. 1979).
@@
In content moderation, research has shown that the accuracy of crowd-sourced hate speech labels differs across groups that are the subjects of hate speech (Davani et al. 2023). For example, hate speech targeting the LGBTQ community was found to be at higher risk of being wrongly labeled as non-hateful (Davani et al. 2023). Such bias in hate speech labels shapes mitigation efforts and thereby undermines the very purpose of safeguarding vulnerable groups.
@@
Despite growing awareness of label bias and its downstream harms, prior work has not proposed scalable and effective methods for identifying mislabeled instances under labeling bias. We refer to this task as bias-aware mislabeling detection. Existing studies that address labeling bias have primarily focused on characterizing label bias and its consequences, or on proposing conceptual strategies to avoid it,such as recommending less biased proxy variables (Obermeyer et al. 2019, Mullainathan and Obermeyer 2021, Passi and Barocas 2019). However, better, bias-proof proxies are often unavailable, or, if they exist, come at a high cost per instance. Critically, extensive evidence shows that many widely used and valuable datasets suffer from entrenched label bias  (Li et al. 2022, Zanger-Tishler et al. 2024, Davani et al. 2023, ProPublica 2016, Obermeyer et al. 2019, Davani et al. 2022, Sap et al. 2019). Effective bias-aware mislabeling detection is thus vital not only for improving the quality of existing datasets but also for guiding label quality assurance in new datasets when bias-free labeling is not feasible to achieve.
@@
Among methodological work, some prior research has aimed to mitigate particular adverse outcomes of label bias with a focus on supervised machine learning tasks (Jiang and Nachum 2020, Wang et al. 2021, Kamiran and Calders 2009). However, such methods do not improve the labels’ quality, but instead focus on improving learning outcomes in the presence of label bias. Importantly, bias-aware mislabeling detection aims to improve the labeling quality at the outset and is thus complementary of such methods, as it aims to improve the reliability of labels in the data itself, which can then be used in a variety of downstream applications. This responds to the common organizational reality in which a single dataset is typically used to support a myriad of data-driven analyses and technologies. As such, our work belongs to the literature on mislabeling detection, where we address a critical gap: existing methodologies assume label error structures that are implicitly incompatible with the presence of label bias.
@@
In this work, we propose and evaluate Decoupled Confident Learning (DeCoLe), a principled machine learning-based framework designed to perform bias-aware mislabeling detection. We first introduce and define the problem of bias-aware mislabeling detection. We then develop DeCoLe, a novel methodology designed to effectively perform bias-aware mislabeling detection. We provide theoretical justification for DeCoLe’s effectiveness in addressing this challenge. We share insights about DeCoLe’s differential performance relative to alternative methods under controlled settings and empirically demonstrate its superior performance in bias-aware mislabeling detection for content moderation—a domain where label bias is a documented challenge. To the best of our knowledge, this is the first empirical evaluation of mislabel detection in the presence of label bias. Crucially, evaluating the effectiveness of mislabeling detection methods is inherently challenging, given that it requires datasets with (1) observed noisy labels, (2) higher-quality gold-standard labels, and (3) demographic group annotations—a combination rarely available in practice due to privacy concerns, organizational liability considerations, and the high cost of collecting high-quality gold-standard labels for large amounts of data. To assess DeCoLe relative to prior mislabeling detection methods, we leverage a rare opportunity afforded by the dataset introduced by Kennedy et al. (2020), which uniquely satisfies these criteria. Empirical results provide robust evidence that DeCoLe substantially outperforms prior methods in bias-aware mislabeling detection for hate speech data. Specifically, DeCoLe identifies more mislabeled instances, especially for the error type that disproportionately affects certain groups, while maintaining high precision of the observed labels in the data estimated as correctly labeled, particularly for the class most impacted by label bias. DeCoLe improves mislabeling detection across all groups without compromising performance for any, challenging the common assumption that correcting bias for some requires sacrificing others and demonstrating that group-aware error detection can benefit all.
@@ 2.1 Data Integrity and Data Quality
Ensuring data integrity and data quality control has been a foundational priority in scientific and applied disciplines for decades. Since the 1970s, scholars across fields—including Information Systems(IS), computer science, and data management—have emphasized the significance of trustworthy data as a prerequisite for reliable decision making and organizational performance (Adams 1975, Ballou and Pazer 1985). These early insights laid the foundation for understanding that high-quality data enables effective decision-making, while poor data quality can undermine outcomes.
@@
In the era of big data—characterized by the 4Vs (Volume, Velocity, Variety, and Veracity) [^1] Big data is characterized by the 4Vs: Volume, denoting its massive scale; Velocity, highlighting the rapid generation and need for timely processing; Variety, reflecting diverse data types; and Veracity, emphasizing the importance of data trustworthiness.(Lu 2020, Cai and Zhu 2015) and propelled by rapidly advancing information technologies (Lu 2020, Cai and Zhu 2015)—, the importance of reliable data to unlock the potential of data-driven insights and analytics is paramount (McKinsey 2014, Davenport et al. 2006). Accordingly, there is an expansion in the number of educational programs (Jafar et al. 2017), designed to equip graduates with the necessary skills to support organizational decision-making, with a strong emphasis on ensuring data integrity as a fundamental priority  (Lu 2020). This reflects a growing awareness that maintaining data integrity is not just a best practice, but a necessity. Poor data quality leads to costly inefficiencies, misinformed decisions, and lost opportunities  (Parssian et al. 2004, Dey and Kumar 2010, Hazen et al. 2014). Ultimately, low-quality data renders information unreliable, undermines the trustworthiness of analytical outputs, compromises usability, and weakens the foundation of decision-support systems (Bai et al. 2012, Agarwal and Dhar 2014, Shmueli 2010). Early detection and correction of errors are therefore essential to prevent distortions from compounding throughout the data pipeline. As errors become more entrenched, the cost of correction rises sharply at each subsequent stage of development  (Currim and Ram 2012), underscoring the need for organizations to proactively assess and address data quality issues at their root (NIST 2023).
@@
Although there is a long-established body of research on data quality, much of it is anchored in the principle of “fitness for use,” emphasizing the role of context in determining data reliability  (Wand and Wang 1996). Data quality is often framed as a multi-dimensional construct, encompassing attributes such as accuracy, completeness, and timeliness. Efforts have been devoted to refining these dimensions and developing methodologies for their measurement (Ballou and Pazer 1985, Krishnan et al. 2005, Dey and Kumar 2010). This stream of work intersects with the “label construct gap” discussed in the next sub-section, which is one of the sources of label bias. However, it is not always clear how to map the macro view of data concepts to the micro usage of data, where each instance carries its own history and story. Existing research largely lacks a systematic approach for bias-aware mislabeling detection in datasets designed for specific tasks, leaving a crucial gap in ensuring high-quality, unbiased data. To address this gap, we propose a principled machine-learning framework specifically designed to detect mislabeled instances in datasets affected by label bias, offering a robust and systematic solution for facilitating data quality improvement at the instance level.
@@ 2.1.1 Mislabeling detection
DeCoLe builds upon a well-established stream of work that focuses on estimating mislabeled instances through principled approaches to noise and error detection (Valiant 1984, Angluin and Laird 1988). Our work is most closely related to the work by Northcutt et al. (2021), which introduces a framework for detecting mislabeled instances via a method we henceforth refer to as Confident Learning (CL) (Northcutt et al. 2021). A foundational aspect of this literature is the classification noise process (CNP) introduced by Angluin and Laird (1988), which assumes that label noise is class-conditional—depending solely on the latent true class. This assumption is reasonable and often appropriate for many datasets  (Goldberger and Ben-Reuven 2017, Sukhbaatar et al. 2014), however, it implicitly rules out the possibility of label bias. In this work, we relax this assumption and propose a methodology for bias-aware mislabeling detection, accounting for the possibility that the likelihood of an erroneous label may depend not only on the latent true class, but also on group membership.

### v2

# Bias-Aware Mislabeling Detection via Decoupled Confident Learning

@@
Abstract
@@ 1   Introduction
In an era defined by rapid digitization and fueled by advances in computational technologies and algorithms, organizations are increasingly reliant on data to drive competitiveness, inform strategy, and support core operations. Evidence from industry underscores the transformative nature of data-driven practices, as data-driven organizations outperform peers by a factor of 23 in customer acquisition, achieve ninefold improvements in customer retention, and realize up to 19 times greater profitability (McKinseyDataMatrics, ; davenport2006competing, ).
@@
Simultaneously, as the strategic value of data-driven practices has intensified, the cost of poor data quality has emerged as a critical point of vulnerability. Analytical systems built on systematic errors in data trigger adverse downstream consequences—producing misleading insights and information that can distort managerial decisions, risk operational disruptions and compliance failures, and erode strategic value and organizational trust (ballou1985modeling, ; bai2012managing, ; krishnan2005data, ; hazen2014data, ; parssian2004assessing, ). The estimated financial implications of poor data quality are profound, with costs reaching 8% to 12% of revenues for typical organizations, potentially leading to billions of dollars in annual losses (hazen2014data, ). Data integrity is at the core of contemporary business competitiveness, which increasingly relies on systematic data auditing and relabeling to enhance data quality (ganju2022electronic, ; liu2024financial, ; bernhardt2022active, ).
@@
A particularly consequential data integrity challenge stems from the quality of labels (frenay2013classification, ). A Label is a covariate that is typically the key focus of an analysis, the result of which often informs impactful decisions and policies at scale and advances knowledge discovery. For instance, a medical diagnosis functions as a label that reflects the presence and severity of a condition and is fundamental to inform treatment decisions. Similarly, in the context of content moderation, a post may be labeled as ‘toxic’ if it contains hate speech or derogatory content (davani2023hate, ), based on which the platform may intervene to mitigate risks. The term “label” has been coined and used extensively in the machine learning literature, while different terms for this concept—such as dependent variable, response variable, outcome, annotation, or rating—exist across disciplines, including social sciences, where these terms similarly refer to an outcome of consequence, central to the motivation and implications of the analyses.
@@
Label bias is a particular form of label error that has attracted significant attention across research communities, including Information Systems (IS) and computer science, because it has been shown to be both common and consequential across domains (obermeyer2019dissecting, ; fu2020artificial, ; bjarnadottir2020machine, ; kokkodis2021demand, ; fu2022fair, ; li2024label, ; jacobs2021measurement, ; passi2019problem, ; suresh2021framework, ; sap2019risk, ; davani2023hate, ; propublica2016machine, ; zanger2024risk, ; hunter1979differential, ; sjoding2020racial, ; fogliato2020fairness, ; akpinar2021effect, ). Label bias refers to systematic errors in the observed labels relative to the gold standard labels, such that the quality of labels differs across social groups (li2022more, ; zanger2024risk, ). Such bias can result in disproportionate harms, with some groups being more adversely affected than others. Specifically, under this bias, the likelihood that an instance is incorrectly labeled depends not only on its true label—a dependency that determines what type of error is more common—but also on its group membership (e.g., age bracket, race, gender, or socioeconomic status). A dependency on both true label and group membership means that the types of errors as well as the rates of errors may vary across groups. Label bias commonly arises in settings where the true label is unobservable or costly to obtain, and observed labels are instead derived from proxies, human judgments or measurement instruments that can embed label bias due to social structural biases, cognitive heuristics, or gaps in domain knowledge (li2022more, ; li2024label, ; passi2019problem, ; jacobs2021measurement, ; eickhoff2018cognitive, ; draws2021checklist, ).
@@
Prior research has revealed the adverse impact of label bias across systems and analyses that utilize the labels, demonstrating that such errors undermine productive data-driven decisions across domains, and potentially perpetuate existing inequalities (li2024label, ; li2022more, ; sap2019risk, ; davani2023hate, ). Analyses that are crucially impacted by labeling quality include data-driven hypothesis generation and pattern discovery, as well as predictive modeling, descriptive modeling, and prescriptive modeling that inform understandings and policies (agarwal2014big, ; shmueli2010explain, ; abbasi2024pathways, ). These risks have also been echoed by federal agencies and policies, including by the National Institute of Standards and Technology (NIST), which urged organizations to enhance data quality  (nist2023ai_rmf, ).
@@
For example, health care systems often aim to estimate health needs to inform the allocation of health resources. However, health needs are often prohibitively costly to estimate. Consequently, health care costs—which are abundant and readily available in the U.S.—have been used as a proxy label for health needs (obermeyer2019dissecting, ; li2022more, ). However, there is a systematic mismatch between health costs and actual health needs, and this mismatch varies across groups. Specifically, because of historical inequities in access to care across patient groups, Black patients have often incurred lower health care costs than white patients with comparable medical conditions. Consequently, a system that relied on cost as a proxy for needs systematically underestimated Black patients’ needs (obermeyer2019dissecting, ). Label quality has been impacted by label bias across a range of other contexts, including criminal justice (propublica2016machine, ; zanger2024risk, ), content moderation davani2023hate ; davani2022dealing ; sap2019risk , and hiring (hunter1979differential, ).
@@
In content moderation, research has shown that the accuracy of crowd-sourced hate speech labels differs across groups that are the subjects of hate speech (davani2023hate, ). For example, hate speech targeting the LGBTQ community was found to be at higher risk of being wrongly labeled as non-hateful (davani2023hate, ). Such bias in hate speech labels shapes mitigation efforts and thereby undermines the very purpose of safeguarding vulnerable groups.
@@
Despite growing awareness of label bias and its downstream harms, prior work has not proposed scalable and effective methods for identifying mislabeled instances under labeling bias. We refer to this task as bias-aware mislabeling detection. Existing studies that address labeling bias have primarily focused on characterizing label bias and its consequences, or on proposing conceptual strategies to avoid it,such as recommending less biased proxy variables (obermeyer2019dissecting, ; mullainathan2021inequity, ; passi2019problem, ). However, better, bias-proof proxies are often unavailable, or, if they exist, come at a high cost per instance. Critically, extensive evidence shows that many widely used and valuable datasets suffer from entrenched label bias  (li2022more, ; zanger2024risk, ; davani2023hate, ; propublica2016machine, ; obermeyer2019dissecting, ; davani2022dealing, ; sap2019risk, ). Effective bias-aware mislabeling detection is thus vital not only for improving the quality of existing datasets but also for guiding label quality assurance in new datasets when bias-free labeling is not feasible to achieve.
@@
Among methodological work, some prior research has aimed to mitigate particular adverse outcomes of label bias with a focus on supervised machine learning tasks (jiang2020identifying, ; wang2021fair, ; kamiran2009classifying, ). However, such methods do not improve the labels’ quality, but instead focus on improving learning outcomes in the presence of label bias. Importantly, bias-aware mislabeling detection aims to improve the labeling quality at the outset and is thus complementary of such methods, as it aims to improve the reliability of labels in the data itself, which can then be used in a variety of downstream applications. This responds to the common organizational reality in which a single dataset is typically used to support a myriad of data-driven analyses and technologies. As such, our work belongs to the literature on mislabeling detection, where we address a critical gap: existing methodologies assume label error structures that are implicitly incompatible with the presence of label bias.
@@
In this work, we propose and evaluate Decoupled Confident Learning (DeCoLe), a principled machine learning-based framework designed to perform bias-aware mislabeling detection. We first introduce and define the problem of bias-aware mislabeling detection. We then develop DeCoLe, a novel methodology designed to effectively perform bias-aware mislabeling detection. We provide theoretical justification for DeCoLe’s effectiveness in addressing this challenge. We share insights about DeCoLe’s differential performance relative to alternative methods under controlled settings and empirically demonstrate its superior performance in bias-aware mislabeling detection for content moderation—a domain where label bias is a documented challenge. To the best of our knowledge, this is the first empirical evaluation of mislabel detection in the presence of label bias. Crucially, evaluating the effectiveness of mislabeling detection methods is inherently challenging, given that it requires datasets with (1) observed noisy labels, (2) higher-quality gold-standard labels, and (3) demographic group annotations—a combination rarely available in practice due to privacy concerns, organizational liability considerations, and the high cost of collecting high-quality gold-standard labels for large amounts of data. To assess DeCoLe relative to prior mislabeling detection methods, we leverage a rare opportunity afforded by the dataset introduced by kennedy2020constructing , which uniquely satisfies these criteria. Empirical results provide robust evidence that DeCoLe substantially outperforms prior methods in bias-aware mislabeling detection for hate speech data. Specifically, DeCoLe identifies more mislabeled instances, especially for the error type that disproportionately affects certain groups, while maintaining high precision of the observed labels in the data estimated as correctly labeled, particularly for the class most impacted by label bias. DeCoLe improves mislabeling detection across all groups without compromising performance for any, challenging the common assumption that correcting bias for some requires sacrificing others and demonstrating that group-aware error detection can benefit all.
@@ 2.1   Data Integrity and Data Quality
Ensuring data integrity and data quality control has been a foundational priority in scientific and applied disciplines for decades. Since the 1970s, scholars across fields—including Information Systems(IS), computer science, and data management—have emphasized the significance of trustworthy data as a prerequisite for reliable decision making and organizational performance (adams1975management, ; ballou1985modeling, ). These early insights laid the foundation for understanding that high-quality data enables effective decision-making, while poor data quality can undermine outcomes.
@@
In the era of big data—characterized by the 4Vs (Volume, Velocity, Variety, and Veracity) [^1] Big data is characterized by the 4Vs: Volume, denoting its massive scale; Velocity, highlighting the rapid generation and need for timely processing; Variety, reflecting diverse data types; and Veracity, emphasizing the importance of data trustworthiness.(lu2020data, ; cai2015challenges, ) and propelled by rapidly advancing information technologies (lu2020data, ; cai2015challenges, )—, the importance of reliable data to unlock the potential of data-driven insights and analytics is paramount (McKinseyDataMatrics, ; davenport2006competing, ). Accordingly, there is an expansion in the number of educational programs (jafar2017emergence, ), designed to equip graduates with the necessary skills to support organizational decision-making, with a strong emphasis on ensuring data integrity as a fundamental priority  (lu2020data, ). This reflects a growing awareness that maintaining data integrity is not just a best practice, but a necessity. Poor data quality leads to costly inefficiencies, misinformed decisions, and lost opportunities  (parssian2004assessing, ; dey2010reassessing, ; hazen2014data, ). Ultimately, low-quality data renders information unreliable, undermines the trustworthiness of analytical outputs, compromises usability, and weakens the foundation of decision-support systems (bai2012managing, ; agarwal2014big, ; shmueli2010explain, ). Early detection and correction of errors are therefore essential to prevent distortions from compounding throughout the data pipeline. As errors become more entrenched, the cost of correction rises sharply at each subsequent stage of development  (currim2012modeling, ), underscoring the need for organizations to proactively assess and address data quality issues at their root (nist2023ai_rmf, ).
@@
Although there is a long-established body of research on data quality, much of it is anchored in the principle of “fitness for use,” emphasizing the role of context in determining data reliability  (wand1996anchoring, ). Data quality is often framed as a multi-dimensional construct, encompassing attributes such as accuracy, completeness, and timeliness. Efforts have been devoted to refining these dimensions and developing methodologies for their measurement (ballou1985modeling, ; krishnan2005data, ; dey2010reassessing, ). This stream of work intersects with the “label construct gap” discussed in the next sub-section, which is one of the sources of label bias. However, it is not always clear how to map the macro view of data concepts to the micro usage of data, where each instance carries its own history and story. Existing research largely lacks a systematic approach for bias-aware mislabeling detection in datasets designed for specific tasks, leaving a crucial gap in ensuring high-quality, unbiased data. To address this gap, we propose a principled machine-learning framework specifically designed to detect mislabeled instances in datasets affected by label bias, offering a robust and systematic solution for facilitating data quality improvement at the instance level.
@@ 2.1.1   Mislabeling detection
DeCoLe builds upon a well-established stream of work that focuses on estimating mislabeled instances through principled approaches to noise and error detection (valiant1984theory, ; angluin1988learning, ). Our work is most closely related to the work by northcutt2021confident , which introduces a framework for detecting mislabeled instances via a method we henceforth refer to as Confident Learning (CL) (northcutt2021confident, ). A foundational aspect of this literature is the classification noise process (CNP) introduced by angluin1988learning , which assumes that label noise is class-conditional—depending solely on the latent true class. This assumption is reasonable and often appropriate for many datasets  (goldberger2017training, ; sukhbaatar2014training, ), however, it implicitly rules out the possibility of label bias. In this work, we relax this assumption and propose a methodology for bias-aware mislabeling detection, accounting for the possibility that the likelihood of an erroneous label may depend not only on the latent true class, but also on group membership.